Write the balanced dissociation reactions for each of the two starting salts. Do not forget to include appropriate charges on ions and indicate phases. Circle ions that are found in the insoluble precipitate you will be making in this lab. Give the names and formula of the precipitate.

2. 0.349 g of CaCl2•2H2O and 0.698 g of Na2CO3 are dissolved in 100 mL of water to form a solution. a. What is the limiting reactant? b. How many grams of CaCO3 will precipitate? c. How many grams of the excess reactant remains after the reaction has gone to completion?

3. This lab will require you to make and then collect a precipitate. Explain how the precipitate will be collected and how its mass will be determined.
Answer #1

The dissociation reactions are:

CaCl2 = Ca + 2 + 2 Cl-

Na2CO3 = 2 Na + + CO3-2

The ions present in the precipitate are: Ca + 2 and CO3-2

a) The moles of the reagents are calculated:

n CaCl2.2H2O = g / MM = 0.349 g / 147 g / mol = 0.0024 mol

n Na2CO3 = 0.698 g / 106 g / mol = 0.0066 mol

The reaction that occurs is:

CaCl2 + Na2CO3 = CaCO3 + 2 NaCl

The limit reagent is CaCl2.

b) The mass of CaCO3 that precipitates is calculated:

m CaCO3 = 0.0024 mol CaCl2 * (1 mol CaCO3 / 1 CaCl2) * (100 g CaCO3 / 1 mol) = 0.24 g

c) The moles and mass of Na2CO3 that remain are calculated:

n Na2CO3 re = 0.0066 - 0.0024 = 0.0042 mol

m Na2CO3 re = 0.0042 mol * 106 g / mol = 0.445 g

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.